# Algorithms with Calibrated Machine Learning Predictions

### v2

@@ Abstract
The field of algorithms with predictions incorporates machine learning advice in the design of online algorithms to improve real-world performance. While this theoretical framework often assumes uniform reliability across all predictions, modern machine learning models can now provide instance-level uncertainty estimates. In this paper, we propose calibration as a principled and practical tool to bridge this gap, demonstrating the benefits of calibrated advice through two case studies: the ski rental and online job scheduling problems. For ski rental, we design an algorithm that achieves optimal prediction-dependent performance and prove that, in high-variance settings, calibrated advice offers more effective guidance than alternative methods for uncertainty quantification. For job scheduling, we demonstrate that using a calibrated predictor leads to significant performance improvements over existing methods. Evaluations on real-world data validate our theoretical findings, highlighting the practical impact of calibration for algorithms with predictions.
@@ 1 Introduction
In recent years, advances in machine learning (ML) models have inspired researchers to revisit the design of classic online algorithms, incorporating insights from ML-based advice to improve decision-making in real-world environments. This research area, termed algorithms with predictions, seeks to design algorithms that are both robust to worst-case inputs and achieve performance that improves with prediction accuracy (a desideratum termed consistency) (Lykouris and Vassilvitskii, 2018). Many prediction-aided algorithms have been developed for online decision-making tasks ranging from rent-or-buy problems like ski rental (Purohit et al., 2018, Anand et al., 2020, Sun et al., 2024) to sequencing problems like job scheduling (Cho et al., 2022).
Algorithms in this framework often rely on global uncertainty parameters intended to summarize the trustworthiness of all of the model’s predictions, with extreme settings indicating that predictions are either all perfect or all uninformative (e.g., Mahdian et al., 2007, Lykouris and Vassilvitskii, 2018, Purohit et al., 2018, Rohatgi, 2020, Wei and Zhang, 2020, Antoniadis et al., 2020). However, ML models often produce local, input-specific uncertainty estimates, exposing a disconnect between theory and practice. For instance, many neural networks provide calibrated probabilities or confidence intervals for each data point. In this paper, we demonstrate that calibration can serve as a powerful tool to bridge this gap. An ML predictor is said to be calibrated if the probabilities it assigns to events match their observed frequencies; when the model outputs a high probability, the event is indeed likely, and when it assigns a low probability, the event rarely occurs. Calibrated predictors convey their uncertainty on each prediction, allowing decision-makers to safely rely on the model’s advice, and eliminating the need for global uncertainty estimates. Moreover, calibrating models after training can be easily achieved through popular methods (e.g. Platt Scaling (Platt et al., 1999) or Histogram Binning (Zadrozny and Elkan, 2001)) and reduces overconfidence, particularly in neural network models (Vasilev and D’yakonov, 2023).
Although we are the first to study calibration for algorithms with predictions, Sun et al. (2024) proposed using conformal prediction in this setting—a common tool in uncertainty quantification (Vovk et al., 2005, Shafer and Vovk, 2008). Conformal predictions provide instance-specific confidence intervals that cover the true label with high probability. We prove that calibration offers key advantages over conformal prediction, especially when the predicted quantities have high variance. In extreme cases, conformal intervals can become too wide to be informative: for binary predictions, a conformal approach returns $\{0,1\}$ unless the true label is nearly certain to be 00 or $1$. In contrast, calibration still conveys information that aids decision-making.
@@ 1.1 Our contributions
The ski rental problem serves as a prototypical example of a broad family of online rent-or-buy problems, where one must choose between an inexpensive, short-term option (renting) and a more costly, long-term option (buying). In this problem, a skier will ski for an unknown number of days and, each day, must decide to either rent skis or pay a one-time cost to buy them. Generalizations of the ski rental problem have informed a broad array of practical applications in networking (Karlin et al., 2001), caching (Karlin et al., 1988), and cloud computing (Khanafer et al., 2013). We design an online algorithm for ski rental that incorporates predictions from a calibrated predictor. We prove that our algorithm achieves optimal expected prediction-level performance for general distributions over instances and calibrated predictors. At a distribution level, its performance degrades smoothly as a function of the mean-squared error and calibration error of the predictor. Moreover, we demonstrate that calibrated predictions can be more informative than the conformal predictions of Sun et al. (2024) when the distribution over instances has high variance that is not explained by features, leading to better performance.
@@
We next study online scheduling in a setting where each job has an urgency level, but only a machine-learned estimate of that urgency is available. This framework is motivated by scenarios such as medical diagnostics, where machine-learning tools can flag potentially urgent cases but cannot fully replace human experts. We demonstrate that using a calibrated predictor provides significantly better guarantees than prior work (Cho et al., 2022), which approached this problem by ordering jobs based on the outputs of a binary predictor. We identify that this method implicitly relies on a crude form of calibration that assigns only two distinct values, resulting in many ties that must be broken randomly. In contrast, we prove that a properly calibrated predictor with finer-grained confidence levels provides a more nuanced job ordering, rigorously quantifying the resulting performance gains.
@@ 1.2 Related work
There has been significant recent interest in integrating ML advice into the design of online algorithms (see, e.g., Mitzenmacher and Vassilvitskii (2022) for a survey). Much of the research in this area assumes uniform uncertainty over prediction reliability (e.g., Lykouris and Vassilvitskii, 2018, Purohit et al., 2018, Wei and Zhang, 2020). Subsequent work has studied more practical settings, assuming access to ML predictors learned from samples (Anand et al., 2020), with probabilistic correctness guarantees (Gupta and Ramdas, 2021), or in the case of binary predictions, with a known confusion matrix (Cho et al., 2022). Recently, Sun et al. (2024) proposed a framework for quantifying prediction-level uncertainty based on conformal prediction. We prove that calibration provides key advantages over conformal prediction in this context, particularly when predicted quantities exhibit high variance.
@@ 2 Preliminaries
Prediction-aided algorithm design.
@@
If $\textnormal{ALG}(\mathcal{A},v,i)$ is the cost incurred by algorithm $\mathcal{A}$ with prediction $f(X)=v$ on instance $i\in\mathcal{I}$, and $\textnormal{OPT}(i)$ is that of the offline optimal solution, the goal is to minimize the expected competitive ratio (CR) defined multiplicatively as
$\mathop{\mathbb{E}}_{(X,I)\sim\mathcal{D}}\left[\frac{\textnormal{ALG}(\mathcal{A},f(X),I)}{\textnormal{OPT}(I)}\right]$
or additively as
$\mathop{\mathbb{E}}_{(X,I)\sim\mathcal{D}}\left[\textnormal{ALG}(\mathcal{A},f(X),I)-\textnormal{OPT}(I)\right]$
(that is, the performance of $\mathcal{A}$ relative to OPT over $\mathcal{D}$). When $\mathcal{D}$, $f$, and the CR type (multiplicative or additive) are clear from context, we denote the expected competitive ratio as $\mathop{\mathbb{E}}[\textnormal{CR}(\mathcal{A})]$. This measure is consistent with prior work on training predictors from samples for algorithms with predictions (Anand et al., 2020).
@@
When $\mathcal{Y}=\{0,1\}$, the equivalent condition is $\Pr[T(I)=1\mid f(X)]=f(X)$, i.e., $f(X)$ directly represents the probability that $T(I)=1$. Achieving perfect calibration is difficult, so in practice we aim to minimize calibration error, such as the max calibration error, which measures the largest deviation from perfect calibration for any prediction.
@@ 3.1 Setup
A skier plans to ski for an unknown number of days $Z\in\mathbb{N}$ and has two options: buy skis at a one-time cost of $b\in\mathbb{N}$ dollars or rent them for $1$ dollar per day. The goal is to determine how many days to rent before buying, minimizing the total cost. If $Z=z$ were known a priori, the optimal policy would rent for $b$ days when $z<b$ and buy immediately otherwise, costing $\min\{z,b\}$. Without knowledge of $z$, competitive ratios of 2 (Karlin et al., 1988) and $\frac{e}{e-1}$ (Karlin et al., 1994) are tight for deterministic and random strategies, respectively. For convenience, we study a continuous variant of this problem where $Z,b,k\in\mathbb{R}_{\geq 0}$ as in prior work (Anand et al., 2020, Sun et al., 2024).
@@
Prediction-aided ski rental.
A deterministic prediction-aided algorithm $\mathcal{A}_{k}$ for ski rental takes as input a prediction $f(X)=v$ and returns a recommendation: “rent skis for $k(v)$ days before buying.” The cost of following this policy when skiing for $z$ days is
$\textnormal{ALG}(\mathcal{A}_{k},v,z)=\begin{cases}k(v)+b&\text{if $z>k(v)$}\\ z&\text{if $z\leq k(v)$}\end{cases}.$
Our goal is to select $k:[0,1]\to\mathbb{R}_{+}$ that minimizes the multiplicative expected CR, denoted $\mathop{\mathbb{E}}[\textnormal{CR}(\mathcal{A}_{k})]$.
@@ 3.2 Ski rental with calibrated predictions
In Algorithm 1, we introduce a deterministic policy for ski rental based on calibrated predictions. To avoid following bad advice, the algorithm defaults to a worst-case strategy of renting for $b$ days unless the prediction is confident that the skier will ski for at least $b$ days. In this second case, the algorithm smoothly interpolates between a strategy that rents for $b$ days and one that rents for $b\sqrt{\alpha/(1+\alpha)}$ days, where $\alpha\in[0,1]$ is a bound on local calibration error that hedges against greedily following predictions.
@@
Condition $\textnormal{OPT}(z)$ $\textnormal{ALG}(\mathcal{A}_{k},v,z)$
@@
(1) and (2) follow from the fact that $f$ predicts $\mathbbm{1}_{C}$ with max calibration error $\alpha$. Under $C=\{Z\geq b\}$, one of conditions (iii) or (iv) from Table 1 hold. In either case, $\textnormal{ALG}(\mathcal{A}_{k},v,Z)/\textnormal{OPT}(Z)\leq 1+\frac{k(v)}{b}$. Under $C^{c}$, one of conditions (i) or (ii) hold. $\textnormal{CR}(\mathcal{A}_{k})=1$ for (i). For (ii),
$\frac{\textnormal{ALG}(\mathcal{A}_{k},v,Z)}{\textnormal{OPT}(Z)}\leq\frac{k(v)+b}{k(v)}=1+\frac{b\cdot\mathbbm{1}_{\{k(v)<b\}}}{k(v)}.$
@@
$\mathop{\mathbb{E}}[\textnormal{CR}(\mathcal{A}_{k})\mid f(X)=v]\leq 1+2\alpha+\frac{(v+\alpha)k(v)}{b}$ $+\mathbbm{1}_{\{k(v)<b\}}\cdot\frac{(1-v+\alpha)b}{k(v)}.$ (2)
@@
Moreover, no deterministic prediction-aided algorithm for ski rental can outperform Algorithm 1 for general distributions $\mathcal{D}$ and calibrated predictors $f$. The construction is non-trivial, so we refer the reader to the proof in Appendix A.
@@
By the tower property of conditional expectation,
$\mathop{\mathbb{E}}[\textnormal{CR}(\mathcal{A}_{k_{*}})]=\mathop{\mathbb{E}}\bigl{[}\mathop{\mathbb{E}}[\textnormal{CR}(\mathcal{A}_{k_{*}})\mid f(X)]\bigr{]}.$
Applying Theorem 3.3 yields
$\mathop{\mathbb{E}}[\textnormal{CR}(\mathcal{A}_{k_{*}})]\leq 1+2\alpha+\mathop{\mathbb{E}}\biggl{[}\min\bigl{\{}f(X)+\alpha,2\sqrt{(f(X)+\alpha)(1-f(X)+\alpha)}\bigr{\}}\biggr{]}.$
@@
$\quad\quad\quad\min\bigl{\{}\mathop{\mathbb{E}}[f(X)]+\alpha,2\sqrt{\mathop{\mathbb{E}}[(f(X)+\alpha)(1-f(X)+\alpha)]}\bigr{\}}.$
@@ 3.3 Comparison to previous work
It is well known that for $\lambda\in(0,1)$, any $(1+\lambda)$-consistent algorithm for deterministic ski rental must be at least $(1+\frac{1}{\lambda})$-robust (Wei and Zhang, 2020, Angelopoulos et al., 2020, Gollapudi and Panigrahi, 2019). While Algorithm 1 is subject to this trade-off in the worst case, calibration provides sufficient information to hedge against adversarial inputs in expectation, leading to substantial improvements in average-case performance. Indeed, it can be seen from the bound in Theorem 3.3 that Algorithm 1 is 1-consistent and always satisfies $\mathop{\mathbb{E}}[\textnormal{CR}(\mathcal{A}_{k_{*}})]\leq 1.8$ when advice is calibrated.
Algorithm 2 Sun et al. (2024) Optimal ski rental with conformal predictions
@@
We are not the first to explore uncertainty quantified predictions for ski rental. Sun et al. (2024) take an orthogonal approach based on conformal prediction. Their method, Algorithm 2, assumes access to a probabilistic interval predictor $\textsc{PIP}_{\delta}:\mathcal{X}\to\mathcal{P}([0,1])$. $\textsc{PIP}_{\delta}$ outputs an interval $[\ell,u]=\textsc{PIP}_{\delta}(X)$ containing the true number of days skied $Z\in[\ell,u]$ with probability at least $1-\delta$. Interval predictions are especially useful when the uncertainty $\delta$ and returned interval width $u-\ell$ are both small. However, as features become less informative, the width of prediction intervals must increase to maintain the same confidence level. This can result in intervals that are too wide to provide meaningful insight into the true number of days skied. Lemma 3.6 and Theorem 3.7 demonstrate that there are infinite families of distributions for which calibrated predictions are more informative than conformal predictions for ski rental.
@@
The construction places mass $a$ on some day $z_{1}\geq 2b$ and mass $1-a$ on $z_{2}\leq\frac{b}{2}$. Any $\textsc{PIP}_{\delta}$ with $\delta<a$ must output an interval $[\ell,u]$ containing both $z_{1}$ and $z_{2}$. Moreover, $\zeta(\delta,\ell)\geq 2$ and $\delta+\frac{u}{b}\geq 2$ by construction. ∎

### v3

@@ Abstract
The field of algorithms with predictions incorporates machine learning advice in the design of online algorithms to improve real-world performance. A central consideration is the extent to which predictions can be trusted—while existing approaches often require users to specify an aggregate trust level, modern machine learning models can provide estimates of prediction-level uncertainty. In this paper, we propose calibration as a principled and practical tool to bridge this gap, demonstrating the benefits of calibrated advice through two case studies: the ski rental and online job scheduling problems. For ski rental, we design an algorithm that achieves near-optimal prediction-dependent performance and prove that, in high-variance settings, calibrated advice offers more effective guidance than alternative methods for uncertainty quantification. For job scheduling, we demonstrate that using a calibrated predictor leads to significant performance improvements over existing methods. Evaluations on real-world data validate our theoretical findings, highlighting the practical impact of calibration for algorithms with predictions.
Machine Learning, ICML
@@ 1 Introduction
In recent years, advances in machine learning (ML) models have inspired researchers to revisit the design of classic online algorithms, incorporating insights from ML-based advice to improve decision-making in real-world environments. This research area, termed algorithms with predictions, seeks to design algorithms that are both robust to worst-case inputs and achieve performance that improves with prediction accuracy (a desideratum termed consistency) (Lykouris & Vassilvitskii, 2018). Many learning-augmented algorithms have been developed for online decision-making tasks ranging from rent-or-buy problems like ski rental (Purohit et al., 2018; Anand et al., 2020; Sun et al., 2024) to sequencing problems like job scheduling (Cho et al., 2022).
This framework often produces a family of algorithms indexed by a single parameter intended to reflect the global reliability of the ML advice. Extreme settings of this parameter yield algorithms that make decisions as if the predictions are either all perfect or all uninformative (e.g., Mahdian et al., 2007; Lykouris & Vassilvitskii, 2018; Purohit et al., 2018; Rohatgi, 2020; Wei & Zhang, 2020; Antoniadis et al., 2020). In contrast, ML models often produce local, prediction-specific uncertainty estimates, exposing a disconnect between theory and practice. For instance, many neural networks provide calibrated probabilities or confidence intervals for each data point.
In this paper, we demonstrate that calibration can serve as a powerful tool to bridge this gap. An ML predictor is said to be calibrated if the probabilities it assigns to events match their observed frequencies; when the model outputs a high probability, the event is indeed likely, and when it assigns a low probability, the event rarely occurs. Calibrated predictors convey their uncertainty on each prediction, allowing decision-makers to safely rely on the model’s advice, and eliminating the need for ad-hoc reliability estimates. Moreover, calibrating an ML model can easily be accomplished using popular methods (e.g. Platt Scaling (Platt et al., 1999) or Histogram Binning (Zadrozny & Elkan, 2001)) that reduce overconfidence (Vasilev & D’yakonov, 2023).
Although we are the first to study calibration for algorithms with predictions, Sun et al. (2024) proposed using conformal prediction in this setting—a common tool in uncertainty quantification (Vovk et al., 2005; Shafer & Vovk, 2008). Conformal predictions provide instance-specific confidence intervals that cover the target with high probability. While these approaches are orthogonal, we prove that calibration can offer key advantages over conformal prediction, especially when the predicted quantities have high variance. In extreme cases, conformal intervals can become too wide to be informative: for binary predictions, a conformal approach returns $\{0,1\}$ unless the true label is nearly certain to be 00 or $1$. In contrast, calibration still conveys information that aids decision-making.
@@ 1.1 Our contributions
The ski rental problem serves as a prototypical example of a broad family of online rent-or-buy problems, where one must choose between an inexpensive, short-term option (renting) and a more costly, long-term option (buying). In this problem, a skier will ski for an unknown number of days and, each day, must decide to either rent skis or pay a one-time cost to buy them. Generalizations of the ski rental problem have informed a broad array of practical applications in networking (Karlin et al., 2001), caching (Karlin et al., 1988), and cloud computing (Khanafer et al., 2013).
We design an online algorithm for ski rental that incorporates predictions from a calibrated predictor. We prove that our algorithm achieves optimal expected prediction-level performance for general distributions over instances and calibrated predictors. At a distribution level, its performance degrades smoothly as a function of the mean-squared error and calibration error of the predictor. Moreover, we demonstrate that calibrated predictions can be more informative than the conformal predictions of Sun et al. (2024) when the distribution over instances has high variance that is not explained by features, leading to better performance.
@@
We next study online scheduling in a setting where each job has an urgency level, but only a machine-learned estimate of that urgency is available. This framework is motivated by scenarios such as medical diagnostics, where machine-learning tools can flag potentially urgent cases but cannot fully replace human experts.
We demonstrate that using a calibrated predictor provides significantly better guarantees than prior work (Cho et al., 2022), which approached this problem by ordering jobs based on the outputs of a binary predictor. We identify that this method implicitly relies on a crude form of calibration that assigns only two distinct values, resulting in many ties that must be broken randomly. In contrast, we prove that a properly calibrated predictor with finer-grained confidence levels provides a more nuanced job ordering, rigorously quantifying the resulting performance gains.
@@ 1.2 Related work
There has been significant recent interest in integrating ML advice into the design of online algorithms (see, e.g., Mitzenmacher & Vassilvitskii (2022) for a survey). Much of the research provides a parameterized family of algorithms with no assumption on the reliability of predictions (e.g., Lykouris & Vassilvitskii, 2018; Purohit et al., 2018; Wei & Zhang, 2020). Subsequent work has studied more practical settings, such as assuming access to ML predictors learned from samples (Anand et al., 2020), with probabilistic correctness guarantees (Gupta et al., 2022), with a known confusion matrix (Cho et al., 2022), or that provide distributional predictions (Dinitz et al., 2024; Angelopoulos et al., 2024; Lin et al., 2022; Diakonikolas et al., 2021). While conceptually related, these papers do not study uncertainty quantification.
Recently, Sun et al. (2024) proposed a framework for quantifying prediction-level uncertainty based on conformal prediction. We show that calibration can offer key advantages over conformal prediction in this context, particularly when predicted quantities exhibit high variance.
@@ 2 Preliminaries
Learning-augmented algorithm design.
@@
If $\mathcal{A}(v,i)$ is the cost incurred by algorithm $\mathcal{A}$ with prediction $f(X)=v$ on instance $i\in\mathcal{I}$, and $\textnormal{OPT}(i)$ is that of the offline optimal solution, the goal is to minimize either the expected competitive ratio (CR)
$\mathop{\mathbb{E}}_{(X,I)\sim\mathcal{D}}\left[\frac{\mathcal{A}(f(X),I)}{\textnormal{OPT}(I)}\right]$
or the expected additive regret $\mathop{\mathbb{E}}\left[\mathcal{A}(f(X),I)-\textnormal{OPT}(I)\right]$, depending on context. Both measure the performance of $\mathcal{A}$ relative to OPT over $\mathcal{D}$. The former is consistent with prior work on training predictors from samples for algorithms with predictions (Anand et al., 2020), while the latter is commonly used to quantify suboptimality in learning-augmented scheduling (Lindermayr & Megow, 2022; Im et al., 2023). When $\mathcal{D}$ and $f$ are clear from context, we refer to these quantities as $\mathop{\mathbb{E}}[\textnormal{CR}(\mathcal{A})]$ and $\mathop{\mathbb{E}}[\textup{R}(\mathcal{A})]$, respectively.
@@
When $\mathcal{Y}=\{0,1\}$, the equivalent condition $\Pr[T(I)=1\mid f(X)]=f(X)$ requires that $f(X)$ is a reliable probabilistic estimate of the event $\{T(I)=1\}$.
A classic result from the literature on probabilistic forecasting states that calibrated predictions are the global minimizers of proper loss functions (DeGroot & Fienberg, 1983). However, achieving perfect calibration is difficult in practice. As a result, post-hoc calibration methods aim to minimize calibration error, such as the max calibration error, which measures the largest deviation from perfect calibration for any prediction.
@@
Given any black box ML model and sufficient data, these methods yield a new predictor with a desired level of calibration error with high probability.
@@ 3.1 Setup
A skier plans to ski for an unknown number of days $Z\in\mathbb{N}$ and has two options: buy skis at a one-time cost of $b\in\mathbb{N}$ dollars or rent them for $1$ dollar per day. The goal is to determine how many days to rent before buying, minimizing the total cost. If $Z=z$ were known a priori, the optimal policy would rent for $b$ days when $z<b$ and buy immediately otherwise, costing $\min\{z,b\}$. Without knowledge of $z$, competitive ratios of 2 (Karlin et al., 1988) and $\frac{e}{e-1}$ (Karlin et al., 1994) are tight for deterministic and random strategies, respectively. For convenience, we study a continuous variant of this problem where $Z,b,k\in\mathbb{R}_{\geq 0}$ as in prior work (Anand et al., 2020; Sun et al., 2024).
@@
Learning-augmented ski rental.
A deterministic learning-augmented algorithm $\mathcal{A}_{k}$ for ski rental takes as input a prediction $f(X)=v$ and returns a recommendation: “rent skis for $k(v)$ days before buying.” The cost of following this policy when skiing for $z$ days is
$\mathcal{A}_{k}(v,z)=\begin{cases}k(v)+b&\text{if $z>k(v)$}\\ z&\text{if $z\leq k(v)$}\end{cases}.$
We aim to select $k:[0,1]\to\mathbb{R}_{+}$ to minimize $\mathop{\mathbb{E}}[\textnormal{CR}(\mathcal{A}_{k})]$.
@@ 3.2 Ski rental with calibrated predictions
In Algorithm 1, we introduce a deterministic policy for ski rental based on calibrated predictions. To avoid following bad advice, the algorithm defaults to a worst-case strategy of renting for $b$ days unless sufficiently confident that the skier will ski for at least $b$ days. In this second case, the algorithm smoothly interpolates between a strategy that rents for $b\sqrt{(1-\alpha)/\alpha}$ days and one that rents for $b\sqrt{\alpha/(1+\alpha)}$ days, where $\alpha\in[0,1]$ is a bound on local calibration error that hedges against greedily following predictions.
@@
Condition $\textnormal{OPT}(z)$ $\mathcal{A}_{k}(v,z)$
@@
(1) and (2) follow from the fact that $f$ predicts $\mathbbm{1}_{C}$ with max calibration error $\alpha$. Under $C=\{Z\geq b\}$, one of conditions (iii) or (iv) from Table 1 hold. In either case, $\mathcal{A}_{k}(v,Z)/\textnormal{OPT}(Z)\leq 1+\frac{k(v)}{b}$. Under $C^{c}$, one of conditions (i) or (ii) hold. $\textnormal{CR}(\mathcal{A}_{k})=1$ for (i). For (ii),
$\frac{\mathcal{A}_{k}(v,Z)}{\textnormal{OPT}(Z)}\leq\frac{k(v)+b}{k(v)}=1+\frac{b\cdot\mathbbm{1}_{\{k(v)<b\}}}{k(v)}.$
@@
$\mathop{\mathbb{E}}[\textnormal{CR}(\mathcal{A}_{k})\mid f(X)=v]\leq$ (2)
$1+2\alpha+\frac{(v+\alpha)k(v)}{b}$ $+\mathbbm{1}_{\{k(v)<b\}}\cdot\frac{(1-v+\alpha)b}{k(v)}.$
@@
$\mathop{\mathbb{E}}[\textnormal{CR}(\mathcal{A}_{k_{*}})\mid f(X)=v]$$\leq$
$1+2\alpha+\min\bigl{\{}v+\alpha$ $,2\sqrt{(v+\alpha)(1-v+\alpha)}\bigr{\}}.$
@@
Moreover, no deterministic learning-augmented algorithm for ski rental can outperform Algorithm 1 for general distributions $\mathcal{D}$ and calibrated predictors $f$. The construction is non-trivial, so we refer the reader to the proof in Appendix A.
@@
By the tower property of conditional expectation, $\mathop{\mathbb{E}}[\textnormal{CR}(\mathcal{A}_{k_{*}})]=\mathop{\mathbb{E}}\bigl{[}\mathop{\mathbb{E}}[\textnormal{CR}(\mathcal{A}_{k_{*}})\mid f(X)]\bigr{]}$. Applying Theorem 3.3 yields
$\mathop{\mathbb{E}}[\textnormal{CR}(\mathcal{A}_{k_{*}})]\leq 1+2\alpha$
$+$ $\mathop{\mathbb{E}}\biggl{[}\min\bigl{\{}f(X)+\alpha,2\sqrt{(f(X)+\alpha)(1-f(X)+\alpha)}\bigr{\}}\biggr{]}.$
@@
$\min\bigl{\{}\mathop{\mathbb{E}}[f(X)]+\alpha,2\sqrt{\mathop{\mathbb{E}}[(f(X)+\alpha)(1-f(X)+\alpha)]}\bigr{\}}.$
@@ 3.3 Comparison to previous work
Consistency and robustness.
It is well known that for $\lambda\in(0,1)$, any $(1+\lambda)$-consistent algorithm for deterministic ski rental must be at least $(1+\frac{1}{\lambda})$-robust (Wei & Zhang, 2020; Angelopoulos et al., 2020; Gollapudi & Panigrahi, 2019). While Algorithm 1 is subject to this trade-off in the worst case, calibration provides sufficient information to hedge against adversarial inputs in expectation, leading to substantial improvements in average-case performance. Indeed, it can be seen from the bound in Theorem 3.3 that Algorithm 1 is 1-consistent and always satisfies $\mathop{\mathbb{E}}[\textnormal{CR}(\mathcal{A}_{k_{*}})]\leq 1.8$ when advice is calibrated ($\alpha=0$). An analysis similar to that of Theorem 15 in Anand et al. (2020) shows that Algorithm 1 is $g(\alpha)$-robust, where
$g(\alpha)=\begin{cases}1+\sqrt{\frac{1+\alpha}{\alpha}}&\text{if $\alpha<1/3$}\\ 2&\text{if $\alpha\geq 1/3$}\end{cases}$
is a decreasing function of $\alpha$. This is because Algorithm 1 executes a worst-case 2-competitive strategy when $\alpha\geq 1/3$ and never buys skis before day $b\sqrt{\frac{\alpha}{1+\alpha}}$ otherwise.
We note that one can run the same algorithm using an artificial upper bound $\alpha^{\prime}>\alpha$ on max calibration error to achieve an improved robustness level $g(\alpha^{\prime})$. As seen from the bounds in Theorem 3.3 and Theorem 3.1, this adjustment will come at the cost of expected performance, highlighting the tradeoff between average and worst-case performance.
Uncertainty quantification.
We are not the first to explore uncertainty quantified predictions for ski rental. Sun et al. (2024) take an orthogonal approach based on conformal prediction. Their method, Algorithm 2, assumes access to a probabilistic interval predictor $\textsc{PIP}_{\delta}:\mathcal{X}\to\mathcal{P}([0,1])$. $\textsc{PIP}_{\delta}$ outputs an interval $[\ell,u]=\textsc{PIP}_{\delta}(X)$ containing the true number of days skied $Z\in[\ell,u]$ with probability at least $1-\delta$. Interval predictions are especially useful when the uncertainty $\delta$ and interval width $u-\ell$ are both small.
However, as features become less informative, the width of prediction intervals must increase to maintain the same confidence level. This can result in intervals that are too wide to provide meaningful insight into the true number of days skied. Lemma 3.6 and Theorem 3.7 demonstrate that there are infinite families of distributions for which calibrated predictions are more informative than conformal predictions for ski rental.
Algorithm 2 (Sun et al., 2024) Optimal ski rental with conformal predictions
@@
The construction places mass $1-a$ on some day $z_{1}\leq\frac{b}{2}$ and mass $a$ on $z_{2}\geq 2b$. Any $\textsc{PIP}_{\delta}$ with $\delta<a$ must output an interval $[\ell,u]$ containing both $z_{1}$ and $z_{2}$. Moreover, $\zeta(\delta,\ell)\geq 2$ and $\delta+\frac{u}{b}\geq 2$ by construction. ∎